Jordan is twice the age of Luke who is 5 years older than Erin. If the total of their ages is 59, how old is Erin?
Answer by edjones(8007)

J+L+E=59
J=2L
E=L-5
.
2L+L+L-5=59 Substitute.
4L-5=59
4L=64
L=16
E=11 answer
J=32
.
Ed